OpenStudy (anonymous):

write the equation in slope intercept form 6x+2y=4 & 6x-6y=12

OpenStudy (the_fizicx99):

Y = mx + b y = -6/2x + 4 y =-6/6x + 12

OpenStudy (anonymous):

No, you're incorrect. $$6x + 2y = 4 \\ 2y = 4 - 6x \\ y = 2 - 3x; \\ 6x - 6y = 12 \\ -6y = 12 - 6x \\ y = -2 + x$$

OpenStudy (anonymous):

The \(b\) term is divided as well.